what is the value of (x-3)(y+6)?
Answer by lenny460(1073)

(x - 3)(y + 6)
Multiply using the FOIL Method.
x * y = xy
x * 6 = 6x
-3 * y = -3y
-3 * 6 = -18
xy + 6x - 3y - 18
The Answer:
xy + 6x - 3y - 18


obuong3@aol.com